A person weighing 75.0 kg (165 lbs) runs a course at 1.78 m/s (4.00 mph). What is the person’s kinetic energy?
1
Expert's answer
2019-08-28T04:26:53-0400

E"_K" = "\frac{1}{2}" (75.0kg)(1.78"\frac{m}{s}") "^2"


= 119 "\frac{ kg.m^2 
2
 }{s^2 
}" = 119J